OpenStudy (anonymous):

Please help! Math test on Friday! y=ln(4x^2-5x)

OpenStudy (anonymous):

\[ \ln(4x^2-5x) = \ln[x(4x-5)] \]

OpenStudy (anonymous):

Use logarithm property: \[ \ln(xy) = \ln(x) + \ln(y) \]

OpenStudy (anonymous):

Or something? What are we trying to do again?